OpenStudy (anonymous):

f(x) = { x-[[x+1]] / (x^2 - 3x + 2), if x<1 { (6-3(sqrt of x)) / (x-4), if x >or=1 Discuss the continuity at x=0, x=1, and x=4

OpenStudy (kc_kennylau):

Just do substitution to see if they are defined :) If the values are defined, it's continuous, otherwise, not.

OpenStudy (anonymous):

Could you show me your answer with solution? :-s

OpenStudy (kc_kennylau):

For example, when x=4, the value is undefined as the denominator equals to 0. Therefore the curve is discontinuous at x=4. Try to do the rest yourself :)

OpenStudy (anonymous):

But when you substitute it in the numerator, you also get 0, meaning you end up with an indeterminate form (0/0). Are you allowed to conjugate when you do f(4)?

OpenStudy (kc_kennylau):

oh sorry then you need to do limit :)

OpenStudy (kc_kennylau):

Ok then do the limit for the 3 cases to see if the limits converge :)

OpenStudy (kc_kennylau):

For example at x=4 the limit is 3/4 which makes it continuous? @kittiwitti1 can you help him :)

OpenStudy (agent0smith):

indeterminate form (0/0) doesn't suddenly make a function continuous. Undefined is undefined.

OpenStudy (kittiwitti1):

@kc_kennylau I don't know if you've read my profile yet but it says I'm still in Pre-Calc...

OpenStudy (agent0smith):

y= sinx/x is not continuous when x=0, even though it's 0/0.

OpenStudy (kc_kennylau):

@kittiwitti1 oh lol sorry

OpenStudy (kittiwitti1):

Haha it's not your fault ^^ I don't blame you :p

OpenStudy (kc_kennylau):

yay :P

OpenStudy (agent0smith):

At x=1, make sure each of the functions on the left and right are equal.

OpenStudy (agent0smith):

Oh you will have to eval. the limit going from the left, and check that equals the function on the right.

OpenStudy (agent0smith):

At x=4 it's just not continuous though.

OpenStudy (anonymous):

But at x=4 when you evaluate f(4) you get -9/12. Then when you get the limit of f(x) as x approaches 4 you also get -9/12 ... Right?

OpenStudy (kc_kennylau):

No, when you evaluate f(4) you get undefined.

OpenStudy (anonymous):

you get (0/0)

OpenStudy (anonymous):

Conjugate f(4) and you get -9/12

OpenStudy (agent0smith):

Why do you keep saying "conjugate"? That's not relevant here.

OpenStudy (agent0smith):

But at x=4 when you evaluate f(4) you get -9/12. Then when you get the limit of f(x) as x approaches 4 you also get -9/12 ... Right? Being continuous means the function MUST be defined AT that point.

OpenStudy (kc_kennylau):

But f(4)=\(\dfrac00\)=undefined\(\dots\)

OpenStudy (anonymous):

0/0 is indeterminate?

OpenStudy (anonymous):

and when you get something indeterminate, you either factor or conjugate?

OpenStudy (kc_kennylau):

undefined is undefined don't try to make it defined lol

OpenStudy (kc_kennylau):

I think you may have confused it with limit where we continuing calculating upon meeting \(\dfrac00\) or \(\dfrac\infty\infty\)

OpenStudy (agent0smith):

Undefined means the function is... undefined at that point. That means discontinuous.

OpenStudy (anonymous):

ohkayyyyy how about x=1?

OpenStudy (agent0smith):

And you're using conjugate very wrongly, i've no idea what you mean by that.

OpenStudy (anonymous):

well..by conjugate i mean multiply the entire thing by 6+3sqr of x / 6 + 3 sqr of x...the numerator ends up 36 - 9x

OpenStudy (anonymous):

oh wait, never mind. forget the x = 4, sorry. i copied the wrong number

OpenStudy (agent0smith):

At x=1, make sure each of the functions on the left and right are equal. Oh you will have to eval. the limit going from the left, and check that equals the function on the right.